# Quantum Calculations of the Cavity Shift in Electron Magnetic Moment Measurements

### v1

@@ 3.2 Cavity and Free Space Results
This yields a shift of the axial level spacing of order $\Delta\omega_{z}\sim(\omega_{z}q/\omega_{c})\Delta\omega_{c}$, where $\Delta\omega_{c}$ is the cavity shift due to Eq. (3.21). By the invariance theorem, this shift yields $\Delta\omega_{c}^{0}\sim\omega_{z}\,\Delta\omega_{z}/\omega_{c}\sim(\omega_{z}^{2}q/\omega_{c}^{2})\Delta\omega_{c}$. Since $q\sim 100$ in the latest iteration of the experiment Fan:2022oyb , this contribution is $\sim 10^{4}$ times smaller than the cyclotron part of the cavity shift, and thus negligible. The contribution from magnetron modes is even more suppressed, since the typical magnetron number is $l\sim 100$ but $\omega_{m}/\omega_{c}\sim 10^{-6}$.
@@ 4 Cavity Shifts in the Relativistic Quantum Theory
However, given that the analogous calculation of the Lamb shift requires a relativistic calculation to match low-energy and high-energy terms, we find it instructive to repeat the calculation in a fully relativistic formalism, taking the nonrelativistic limit only in the final step. Indeed, even the relativistic calculation of the cyclotron shift in free space (which, as discussed in Sec. 2, affects the spacing between cyclotron levels at $\mathcal{O}(B^{3})$) requires a careful matching between low- and high-energy regimes PhysRevD.8.3446 . It is also enlightening to see how the cyclotron shift arises from the same one-loop diagrams that can be used to compute g2g-2.
As we will discuss in Sec. 4.1, the relativistic calculation requires computing the difference of one-loop self-energy diagrams in Fig. 2, involving the in-cavity photon and free-space photon propagators. As for the electron propagator, there is no known exact solution in the full set of trap fields; instead, in Sec. 4.2 we will use Schwinger’s result PhysRev.82.664 for the electron propagator in a constant $B$-field alone, and include the effects of the quadrupole potential through the external-leg electron wavefunctions. As such, here we will not distinguish between $\omega_{c}^{0}$ and $\omega_{c}$.
@@
  • We must perform an additional renormalization to subtract the electron self-energy in the absence of an external magnetic field. This is exactly analogous to mass renormalization in the standard treatment of the Lamb shift itzykson2012quantum , and is absent in the nonrelativistic calculation where the electron mass is not treated as a contribution to the energy.
@@ 4.1 Energy Shifts from Self-Energy Diagrams
First, we review standard arguments (e.g. see Ref. weinberg ) for how self-energy diagrams compute energy-level shifts. The exact position-space propagator for a fermion $\Psi$ is defined as the time-ordered correlation function of fields
@@
In the case where $A_{\mu}$ corresponds to a constant $B$-field, Schwinger derived a non-perturbative expression for $S_{A}$ as an integral over a Schwinger proper time variable, which we will use in detail in Sec. 4.2 below and refer to as the Schwinger propagator PhysRev.82.664 ; kuznetsov2013electroweak . The effect of $\delta S_{A}$ is to shift the energy eigenvalues away from $\pm E_{n}$, the poles of the Schwinger propagator (namely the relativistic Landau levels), and will also shift the spinor eigenstates $u_{n},v_{n}$ of $S_{A}$. Taking $E^{\prime}_{n}=E_{n}+\delta E_{n}$ and $U_{n}=u_{n}+\delta u_{n}$ in Eq. (4.2) and expanding to first order in $\delta$’s, we have
@@
where $S_{A}$ is the electron propagator, $D_{\mu\nu}$ is the photon propagator, and $\tau=t-t^{\prime}$. The two factors of $S_{A}$ in Eq. (4.6) provide the required energy denominators, and taking inner products with a particular unperturbed eigenstate $u_{n}$ to isolate $E_{n}$, the energy shift is given by
@@
The result for $\delta E_{n}$ makes intuitive sense, as it resembles a first-order energy shift in nonrelativistic quantum mechanics, with perturbing “Hamiltonian” given by the amputated one-loop diagram. [^2] At one-loop, there is also a diagram corresponding to vacuum polarization, but as argued in Sec. 2, its effects are negligible here. Indeed, one can check that it vanishes identically in the limit of no quadrupole fields. The setup of our calculation is very similar to Ref. PhysRevD.8.3446 , which used the Schwinger propagator to compute free-space energy shifts to $\mathcal{O}(B^{3})$ for a generic electron state; as a check, we have reproduced those results to $\mathcal{O}(B^{2})$. In addition, Eq. (4.8) can be used to compute the free-space one-loop correction to the spin-flip energy $\omega_{s}$, by taking the states $u_{n}$ to represent spin-down and spin-up electrons with the same cyclotron level index. As shown in Refs. PhysRevD.8.3446 ; Suzuki:2005wra , this recovers the usual one-loop anomalous magnetic moment (g2g-2), more commonly computed by considering the vertex renormalization.
@@ 4.2 Propagators and Wavefunctions
The Schwinger propagator can be written in the form kuznetsov2013electroweak
@@
the Schwinger phase is kuznetsov2013electroweak ; Suzuki:2005wra  [^3] Note that Ref. kuznetsov2013electroweak defines $\Phi$ with an additional minus sign and works in Landau gauge rather than symmetric gauge; by performing a gauge transformation to symmetric gauge, we have checked that our expressions for the wavefunctions and Schwinger phase agree with Ref. Suzuki:2005wra which uses symmetric gauge.
@@
The electron wavefunctions which solve the Dirac equation in the presence of the vector potential $\mathbf{A}$ in Eq. (4.10) are plane waves in $z$. There are no known exact solutions of the Dirac equation once the confining quadrupole potential is included, but the axial confinement of the electron is crucial in order to apply the dipole approximation. As a result, we will take our electron wavefunctions to be the approximate energy eigenstates of the Dirac equation in symmetric gauge kuznetsov2013electroweak ; Suzuki:2005wra ,
@@
where $\bar{z}$ is the $z$-coordinate of the cavity center. By assumption, the electron is localized on the cavity axis at $x=y=0$. We have put in a Gaussian axial confinement, parameterized by $\sigma_{z}$, by hand. However, we will show that the final result is independent of $\sigma_{z}$, indicating that the exact form of the confining wavefunction is irrelevant, as long as the axial confinement is sufficiently weak. The condition we obtain on $\sigma_{z}$ is parametrically identical to the one obtained in Ref. Bordag:1988yd for the parallel-plate geometry.
@@
QED calculations in Coulomb gauge typically also require the use of the longitudinal part of the propagator, $D_{00}$, which arises from the static Coulomb interaction between charges. In our case, because this component of the propagator is time-independent, it will shift all energy levels uniformly and therefore not contribute to the shift of the cyclotron frequency PhysRevD.32.729 . Therefore, we are free to ignore it and use only the transverse propagator $D^{ij}$.
@@ 4.3 Cavity and Free Space Results
The $\sigma_{z}$ dependence cancels in the integrand, $\sqrt{\frac{\sigma_{z}^{2}}{\sigma_{z}^{2}+2is}}\sim 1$, as long as the upper bound of the inequality in (4.35) is much less than 1. For a cavity of size $R$, the lowest-lying cavity mode is of order $\omega_{\sigma s}^{0}\sim 1/R$, so the energy shift is independent of $\sigma_{z}$ as long as $\sigma_{z}\gg\sqrt{R/m}\sim 4\times 10^{-5}\ {\rm mm}$, where we took $R\sim 4\ {\rm mm}$. (This condition was first noted in Bordag:1988yd , where $R$ was the separation between parallel plates.) This is satisfied in typical experimental setups, where $\sigma_{z}\sim 4\times 10^{-4}\ {\rm mm}$.
@@ 5 Spherical Cavity
Now we explicitly compute the cavity shift for a spherical cavity. We express it in terms of the difference of a divergent sum and divergent integral, appropriately regularized, and show that the difference can be evaluated by considering an appropriate contour integral. The final result precisely matches the classical result of Ref. PhysRevA.34.2638 .
@@ 5.1 Setup
The sum is written in terms of a $p^{\text{th}}$ Bessel zero, so to relate it to an integral we must render it well-defined and well-behaved for non-integer pp. [^5] The following technique is due to Olver Olver_1950 , and is further discussed in Ref. 10.1063/1.1665021 ; abramowitz1965handbook . To start, note that the definition of $c_{p}$ is equivalent to $\tan c_{p}=-c_{p}/(c_{p}^{2}-1)$. Inverting this expression gives
@@ 5.3 Discussion
Ref. PhysRevA.34.2638 computed the cavity shift classically; taking its Eq. (12) and plugging in the free-space decay rate (in the cgs units used in that work) gives exactly the same answer.
@@ 6.1 Setup
Next, we rewrite the summand in a way that is well-defined and well-behaved for non-integer $n$. First, for $f_{\mathrm{TM}}(n)$, we start with Olver’s trick (see section 9.5 of Ref. abramowitz1965handbook ), defining
@@ 6.3 Discussion
Figure 5: Numeric values of the cavity shift for the cylinder, computed using our renormalized mode sum result Eq. (6.41), and using the existing classical Green’s function result Eq. (6.46). The two agree perfectly, and cannot be distinguished on the plot. The sums are evaluated up to $n=50$, but summing to $n=10$ already gives a match to $\sim 1\%$ accuracy. The top value of $L/a$ was chosen to match the most recent measurement Fan:2022oyb , and the bottom value was chosen to match that considered in Ref. PhysRevA.32.3204 .
@@
The classical result, given by taking the real part of Eq. (4.23) of Ref. PhysRevA.32.3204 , is
@@
where $\xi=\omega_{c}L$, $k_{p}=(2p+1)\pi/L$ and $\mu_{p}=\sqrt{k_{p}^{2}-\omega_{c}^{2}}$. (We note that here we are taking $L$ to be the cylinder’s total length, but in Ref. PhysRevA.32.3204 it was the half of the cylinder’s length.) The first line of Eq. (6.46) depends only on $L$, and represents the effect of the cavity endcaps. The second line is the correction due to the curved sides of the cavity, and contains poles when the cyclotron frequency is equal to a cavity mode frequency.
@@ 7 Conclusion
The renormalization of the cavity shift is qualitatively similar to that of the Casimir effect, reviewed in Refs. Milton:2004ya ; Jaffe:2005vp ; Lamoreaux:2005gf ; Klimchitskaya:2009cw . Unlike the Lamb shift, which contains an $\mathcal{O}(1)$ high-energy contribution that must be calculated in the relativistic theory, both the cavity shift and Casimir effect can be computed accurately using nonrelativistic electrons. In fact, they both arise far in the infrared, from modes with energies on the scale $1/R$, where $R$ is the length scale of the conductors. Furthermore, we have seen that earlier calculations of the cavity shift using the (self)-field of the electron give the same results as our calculation which directly uses cavity modes. This is reminiscent of how the Casimir effect can be thought of as due to the fields of fluctuating electrons in the cavity walls, as was shown by Lifshitz and Schwinger lifshitz1956theory ; schwinger_1978 , or equivalently as due to the vacuum energy of modes.
On a more technical level, a great deal of formalism has been developed to calculate the Casimir effect in various geometries, which may be of use for future cavity shift calculations. In particular, our contour integration method is similar in spirit to the treatment of the Casimir effect for spherical shells given in Ref. 10.1063/1.1666141 , and it may be possible to apply the generalizations of the Abel–Plana formula derived in Refs. inui2003generalized ; Saharian:2007ph to more general cavity shapes.
@@
As shown in Eq. (2.5), the fractional cavity shifts due to transverse and longitudinal (i.e. electrostatic) fields are proportional to $1/m$ and $m$ respectively. As a result of this scaling, only the transverse field is relevant for trapped electrons, while only the longitudinal field is relevant for trapped ions. In the context of trapped ions, this effect is called the “image charge shift.” It was first measured in 1989 PhysRevA.40.6308 , and is routinely accounted for in precision measurements, e.g. see Refs. vandyck2006231 ; BASE:2014drs ; BASE:2016yuo ; Schneider:2017lff ; PhysRevLett.119.033001 .
Exact solutions for the image charge shift exist in idealized geometries tinkle_2001 ; Winters_2006 , while in cylindrically symmetric geometries a semianalytic solution is available PhysRevA.64.023403 . More generally, it can be determined numerically using the measured cavity geometry to $\sim 1\%$ precision PhysRevA.100.023411 . This approach also applies to the cavities used for trapped electrons, allowing the longitudinal cavity shift to be found with accuracy $\Delta\omega_{c}/\omega_{c}\sim 10^{-16}$, sufficient for the near future.
These results and methods do not seem to be applicable to the transverse cavity shift, which must be separately accounted for. The longitudinal cavity shift is substantially simpler because it can be computed without renormalization, e.g. by considering the Coulomb fields of the static charges induced on the cavity walls. Furthermore, as already noted in Ref. PhysRevA.64.023403 , the transverse cavity shift depends on the cavity’s full mode structure, and is therefore much more sensitive to the details of the cavity geometry and boundary conditions.
@@
Here we will consider the even simpler case of a hard cutoff, evaluating the sum up to $p=p_{0}$ and the integral up to $c=c_{\mathrm{max}}$. This is not a smooth regulator, and one could reasonably take $c_{\mathrm{max}}\in[c_{p_{0}},c_{p_{0}+1}]$. Since $c_{p}\approx\pi p$ for large $p$, this yields fractional cavity shifts that vary by $\sim\alpha/ma$. However, it was shown in Ref. Barton_1981 that for linearly divergent sums, one will obtain the correct answer (i.e. the one obtained for a smooth regulator) if one takes the integral cutoff to lie halfway between sum elements, $c_{\mathrm{max}}=(c_{p_{0}}+c_{p_{0}+1})/2$, as shown in Fig. 6. Furthermore, a close match to the exact answer is obtained even for relatively low $p_{0}$.
@@
The cavity shift has long been one of the dominant systematic uncertainties of electron g2g-2 measurements. In fact, one of the main reasons a cylindrical cavity is used is that the cavity shift in an ideal cylinder has been computed cylinder_1989 .
In practice, the real cavity differs from an ideal cylinder, due to geometrical imperfections and the presence of the trapping electrodes. Thus, in the most recent measurement Fan:2022oyb , several dozen low-lying cavity modes were measured and identified with specific cylindrical cavity mode numbers. Compared to the ideal modes, the measured modes have slightly different frequencies and coupling strengths, as well as finite quality factors. The result for an ideal cylinder is then modified by subtracting out what an ideal mode would have contributed, and adding back the contribution from the corresponding measured mode.

### v2

@@ 3.2 Cavity and Free Space Results
This yields a shift of the axial level spacing of order $\Delta\omega_{z}\sim(\omega_{z}q/\omega_{c})\Delta\omega_{c}$, where $\Delta\omega_{c}$ is the cavity shift due to Eq. (3.21). By the invariance theorem, this shift yields $\Delta\omega_{c}^{0}\sim\omega_{z}\,\Delta\omega_{z}/\omega_{c}\sim(\omega_{z}^{2}q/\omega_{c}^{2})\Delta\omega_{c}$. Since $q\sim 100$ in the latest iteration of the experiment Fan (2022), this contribution is $\sim 10^{4}$ times smaller than the cyclotron part of the cavity shift, and thus negligible. The contribution from magnetron modes is even more suppressed, since the typical magnetron number is $l\sim 100$ but $\omega_{m}/\omega_{c}\sim 10^{-6}$.
@@ 4 Cavity Shifts in the Relativistic Quantum Theory
However, given that the analogous calculation of the Lamb shift requires a relativistic calculation to match low-energy and high-energy terms, we find it instructive to repeat the calculation in a fully relativistic formalism, taking the nonrelativistic limit only in the final step. Indeed, even the relativistic calculation of the cyclotron shift in free space (which, as discussed in Sec. 2, affects the spacing between cyclotron levels at $\mathcal{O}(B^{3})$) requires a careful matching between low- and high-energy regimes Tsai and Yildiz (1973). It is also enlightening to see how the cyclotron shift arises from the same one-loop diagrams that can be used to compute g2g-2.
As we will discuss in Sec. 4.1, the relativistic calculation requires computing the difference of one-loop self-energy diagrams in Fig. 2, involving the in-cavity photon and free-space photon propagators. As for the electron propagator, there is no known exact solution in the full set of trap fields; instead, in Sec. 4.2 we will use Schwinger’s result Schwinger (1951) for the electron propagator in a constant $B$-field alone, and include the effects of the quadrupole potential through the external-leg electron wavefunctions. As such, here we will not distinguish between $\omega_{c}^{0}$ and $\omega_{c}$.
@@
  • We must perform an additional renormalization to subtract the electron self-energy in the absence of an external magnetic field. This is exactly analogous to mass renormalization in the standard treatment of the Lamb shift Itzykson and Zuber (2012), and is absent in the nonrelativistic calculation where the electron mass is not treated as a contribution to the energy.
@@ 4.1 Energy Shifts from Self-Energy Diagrams
First, we review standard arguments (e.g. see Ref. Weinberg (1995)) for how self-energy diagrams compute energy-level shifts. The exact position-space propagator for a fermion $\Psi$ is defined as the time-ordered correlation function of fields
@@
In the case where $A_{\mu}$ corresponds to a constant $B$-field, Schwinger derived a non-perturbative expression for $S_{A}$ as an integral over a Schwinger proper time variable, which we will use in detail in Sec. 4.2 below and refer to as the Schwinger propagator Schwinger (1951); Kuznetsov and Mikheev (2013). The effect of $\delta S_{A}$ is to shift the energy eigenvalues away from $\pm E_{n}$, the poles of the Schwinger propagator (namely the relativistic Landau levels), and will also shift the spinor eigenstates $u_{n},v_{n}$ of $S_{A}$. Taking $E^{\prime}_{n}=E_{n}+\delta E_{n}$ and $U_{n}=u_{n}+\delta u_{n}$ in Eq. (4.2) and expanding to first order in $\delta$’s, we have
@@
where $S_{A}$ is the electron propagator, $D_{\mu\nu}$ is the photon propagator, and $\tau=t-t^{\prime}$. The two factors of $S_{A}$ in Eq. (4.6) provide the required energy denominators, and taking inner products with a particular unperturbed eigenstate $u_{n}$ to isolate $E_{n}$, the energy shift is given by [^3] This procedure has previously been used to calculate the energy level shift due to conducting plates Bordag et al. (1988).
@@
The result for $\delta E_{n}$ makes intuitive sense, as it resembles a first-order energy shift in nonrelativistic quantum mechanics, with perturbing “Hamiltonian” given by the amputated one-loop diagram. [^4] At one-loop, there is also a diagram corresponding to vacuum polarization, but as argued in Sec. 2, its effects are negligible here. Indeed, one can check that it vanishes identically in the limit of no quadrupole fields. The setup of our calculation is very similar to Ref. Tsai and Yildiz (1973), which used the Schwinger propagator to compute free-space energy shifts to $\mathcal{O}(B^{3})$ for a generic electron state; as a check, we have reproduced those results to $\mathcal{O}(B^{2})$. In addition, Eq. (4.8) can be used to compute the free-space one-loop correction to the spin-flip energy $\omega_{s}$, by taking the states $u_{n}$ to represent spin-down and spin-up electrons with the same cyclotron level index. As shown in Refs. Tsai and Yildiz (1973); Suzuki (2005), this recovers the usual one-loop anomalous magnetic moment (g2g-2), more commonly computed by considering the vertex renormalization.
@@ 4.2 Propagators and Wavefunctions
The Schwinger propagator can be written in the form Kuznetsov and Mikheev (2013)
@@
the Schwinger phase is Kuznetsov and Mikheev (2013); Suzuki (2005) [^5] Note that Ref. Kuznetsov and Mikheev (2013) defines $\Phi$ with an additional minus sign and works in Landau gauge rather than symmetric gauge; by performing a gauge transformation to symmetric gauge, we have checked that our expressions for the wavefunctions and Schwinger phase agree with Ref. Suzuki (2005) which uses symmetric gauge.
@@
The electron wavefunctions which solve the Dirac equation in the presence of the vector potential $\mathbf{A}$ in Eq. (4.10) are plane waves in $z$. There are no known exact solutions of the Dirac equation once the confining quadrupole potential is included, but the axial confinement of the electron is crucial in order to apply the dipole approximation. As a result, we will take our electron wavefunctions to be the approximate energy eigenstates of the Dirac equation in symmetric gauge Kuznetsov and Mikheev (2013); Suzuki (2005),
@@
where $\bar{z}$ is the $z$-coordinate of the cavity center. By assumption, the electron is localized on the cavity axis at $x=y=0$. We have put in a Gaussian axial confinement, parameterized by $\sigma_{z}$, by hand. However, we will show that the final result is independent of $\sigma_{z}$, indicating that the exact form of the confining wavefunction is irrelevant, as long as the axial confinement is sufficiently weak. The condition we obtain on $\sigma_{z}$ is parametrically identical to the one obtained in Ref. Bordag et al. (1988) for the parallel-plate geometry.
@@
QED calculations in Coulomb gauge typically also require the use of the longitudinal part of the propagator, $D_{00}$, which arises from the static Coulomb interaction between charges. In our case, because this component of the propagator is time-independent, it will shift all energy levels uniformly and therefore not contribute to the shift of the cyclotron frequency Boulware et al. (1985). Therefore, we are free to ignore it and use only the transverse propagator $D^{ij}$.
@@ 4.3 Cavity and Free Space Results
The $\sigma_{z}$ dependence cancels in the integrand, $\sqrt{\frac{\sigma_{z}^{2}}{\sigma_{z}^{2}+2is}}\sim 1$, as long as the upper bound of the inequality in (4.35) is much less than 1. For a cavity of size $R$, the lowest-lying cavity mode is of order $\omega_{\sigma s}^{0}\sim 1/R$, so the energy shift is independent of $\sigma_{z}$ as long as $\sigma_{z}\gg\sqrt{R/m}\sim 4\times 10^{-5}\ {\rm mm}$, where we took $R\sim 4\ {\rm mm}$. (This condition was first noted in Bordag et al. (1988), where $R$ was the separation between parallel plates.) This is satisfied in typical experimental setups, where $\sigma_{z}\sim 4\times 10^{-4}\ {\rm mm}$.
@@ 5 Spherical Cavity
Now we explicitly compute the cavity shift for a spherical cavity. We express it in terms of the difference of a divergent sum and divergent integral, appropriately regularized, and show that the difference can be evaluated by considering an appropriate contour integral. The final result precisely matches the classical result of Ref. Brown et al. (1986).
@@ 5.1 Setup
The sum is written in terms of a $p^{\text{th}}$ Bessel zero, so to relate it to an integral we must render it well-defined and well-behaved for non-integer pp. [^7] The following technique is due to Olver Olver (1950), and is further discussed in Ref. Boyer (1969); Abramowitz and Stegun (1965). To start, note that the definition of $c_{p}$ is equivalent to $\tan c_{p}=-c_{p}/(c_{p}^{2}-1)$. Inverting this expression gives
@@ 5.3 Discussion
Ref. Brown et al. (1986) computed the cavity shift classically; taking its Eq. (12) and plugging in the free-space decay rate (in the cgs units used in that work) gives exactly the same answer.
@@ 6.1 Setup
Next, we rewrite the summand in a way that is well-defined and well-behaved for non-integer $n$. First, for $f_{\mathrm{TM}}(n)$, we start with Olver’s trick (see section 9.5 of Ref. Abramowitz and Stegun (1965)), defining
@@ 6.3 Discussion
Figure 5: Numeric values of the cavity shift for the cylinder, computed using our renormalized mode sum result Eq. (6.41), and using the existing classical Green’s function result Eq. (6.46). The two agree perfectly, and cannot be distinguished on the plot. The sums are evaluated up to $n=50$, but summing to $n=10$ already gives a match to $\sim 1\%$ accuracy. The top value of $L/a$ was chosen to match the most recent measurement Fan (2022), and the bottom value was chosen to match that considered in Ref. Brown et al. (1985).
@@
The classical result, given by taking the real part of Eq. (4.23) of Ref. Brown et al. (1985), is
@@
where $\xi=\omega_{c}L$, $k_{p}=(2p+1)\pi/L$ and $\mu_{p}=\sqrt{k_{p}^{2}-\omega_{c}^{2}}$. (We note that here we are taking $L$ to be the cylinder’s total length, but in Ref. Brown et al. (1985) it was the half of the cylinder’s length.) The first line of Eq. (6.46) depends only on $L$, and represents the effect of the cavity endcaps. The second line is the correction due to the curved sides of the cavity, and contains poles when the cyclotron frequency is equal to a cavity mode frequency.
@@ 7 Conclusion
The renormalization of the cavity shift is qualitatively similar to that of the Casimir effect, reviewed in Refs. Milton (2004); Jaffe (2005); Lamoreaux (2005); Klimchitskaya et al. (2009). Unlike the Lamb shift, which contains an $\mathcal{O}(1)$ high-energy contribution that must be calculated in the relativistic theory, both the cavity shift and Casimir effect can be computed accurately using nonrelativistic electrons. In fact, they both arise far in the infrared, from modes with energies on the scale $1/R$, where $R$ is the length scale of the conductors. Furthermore, we have seen that earlier calculations of the cavity shift using the (self)-field of the electron give the same results as our calculation which directly uses cavity modes. This is reminiscent of how the Casimir effect can be thought of as due to the fields of fluctuating electrons in the cavity walls, as was shown by Lifshitz and Schwinger Lifshitz (1956); Schwinger et al. (1978), or equivalently as due to the vacuum energy of modes.
On a more technical level, a great deal of formalism has been developed to calculate the Casimir effect in various geometries, which may be of use for future cavity shift calculations. In particular, our contour integration method is similar in spirit to the treatment of the Casimir effect for spherical shells given in Ref. Davies (1972), and it may be possible to apply the generalizations of the Abel–Plana formula derived in Refs. Inui (2003); Saharian (2007) to more general cavity shapes.
Higher-Order Corrections.
Since we have shown that the one-loop self energy gives precisely the same cavity shift as a classical calculation, it is natural to ask whether this agreement continues at higher order. Though we do not have a definitive answer to this question, we can give some heuristic arguments. A two-loop diagram only contributes to the cavity shift in regions where at least one of the loop momenta is at the infrared scale $\sim 1/R$.
First, suppose only one of the loop momenta is at the low scale. By dimensional analysis, one would expect a contribution to the cavity shift scaling as $\Delta\omega_{c}/\omega_{c}\sim\alpha^{2}/(mR)$, which is potentially relevant. Restoring factors of $\hbar$, we have $\alpha\propto 1/\hbar$, and $1/m$ becomes the Compton scale $\hbar/m$, so that this contribution to the cavity shift is not $\hbar$-independent. In this sense, it is not classical. However, from the perspective of the low-energy loop, one can absorb the high-energy loop’s effect into a renormalized mass and charge for the electron. Then our one-loop calculation would already include this effect if one uses the physical electron mass and charge.
There may be a residual contribution from the region where both loop momenta are at the low scale, which would scale roughly as $\Delta\omega_{c}/\omega_{c}\sim(\alpha/mR)^{2}$. This is $\hbar$-independent, and thus could potentially arise from a higher-order classical calculation. In the leading-order classical calculation, one considers the backreaction of the electron’s cavity field on the electron’s motion. At second-order, one could account for how this backreaction modifies the electron’s cavity field, which then acts back on the electron again. We do not know if this contribution would agree with the two-loop calculation, but it will be negligible for the foreseeable future.
@@
As shown in Eq. (2.5), the fractional cavity shifts due to transverse and longitudinal (i.e. electrostatic) fields are proportional to $1/m$ and $m$ respectively. As a result of this scaling, only the transverse field is relevant for trapped electrons, while only the longitudinal field is relevant for trapped ions. In the context of trapped ions, this effect is called the “image charge shift.” It was first measured in 1989 Van Dyck et al. (1989), and is routinely accounted for in precision measurements, e.g. see Refs. Van Dyck et al. (2006); Mooser et al. (2014); Smorra and others (2017); Schneider and others (2017); Heiße et al. (2017).
Exact solutions for the image charge shift exist in idealized geometries Tinkle and Barlow (2001); Winters et al. (2006), while in cylindrically symmetric geometries a semianalytic solution is available Porto (2001). More generally, it can be determined numerically using the measured cavity geometry to $\sim 1\%$ precision Schuh et al. (2019). This approach also applies to the cavities used for trapped electrons, allowing the longitudinal cavity shift to be found with accuracy $\Delta\omega_{c}/\omega_{c}\sim 10^{-16}$, sufficient for the near future.
These results and methods do not seem to be applicable to the transverse cavity shift, which must be separately accounted for. The longitudinal cavity shift is substantially simpler because it can be computed without renormalization, e.g. by considering the Coulomb fields of the static charges induced on the cavity walls. Furthermore, as already noted in Ref. Porto (2001), the transverse cavity shift depends on the cavity’s full mode structure, and is therefore much more sensitive to the details of the cavity geometry and boundary conditions.
@@
Here we will consider the even simpler case of a hard cutoff, evaluating the sum up to $p=p_{0}$ and the integral up to $c=c_{\mathrm{max}}$. This is not a smooth regulator, and one could reasonably take $c_{\mathrm{max}}\in[c_{p_{0}},c_{p_{0}+1}]$. Since $c_{p}\approx\pi p$ for large $p$, this yields fractional cavity shifts that vary by $\sim\alpha/ma$. However, it was shown in Ref. Barton (1981) that for linearly divergent sums, one will obtain the correct answer (i.e. the one obtained for a smooth regulator) if one takes the integral cutoff to lie halfway between sum elements, $c_{\mathrm{max}}=(c_{p_{0}}+c_{p_{0}+1})/2$, as shown in Fig. 6. Furthermore, a close match to the exact answer is obtained even for relatively low $p_{0}$.
@@
The cavity shift has long been one of the dominant systematic uncertainties of electron g2g-2 measurements. In fact, one of the main reasons a cylindrical cavity is used is that the cavity shift in an ideal cylinder has been computed Tan and Gabrielse (1989).
In practice, the real cavity differs from an ideal cylinder, due to geometrical imperfections and the presence of the trapping electrodes. Thus, in the most recent measurement Fan (2022), several dozen low-lying cavity modes were measured and identified with specific cylindrical cavity mode numbers. Compared to the ideal modes, the measured modes have slightly different frequencies and coupling strengths, as well as finite quality factors. The result for an ideal cylinder is then modified by subtracting out what an ideal mode would have contributed, and adding back the contribution from the corresponding measured mode.